OpenStudy (anonymous):

how do i factor x^3+2x^2-5x-6?

jimthompson5910 (jim_thompson5910):

Hint: list out the possible rational roots and use synthetic division and the remainder theorem to find the factors

OpenStudy (anonymous):

ok. i will try to rememeber how to do that. I am iin pre-calc now but need to do this and havent had algebra II for a year and a half!

OpenStudy (nurali):

You can use factor theorem. f(x) = x^3+2x^2-5x-6 f(-1) = -1+2+5-6 = 0 So, x+1 is a factor Use synthetic division, x^3+2x^2-5x-6 = (x+1)(x^2+x-6) = (x+1)(x-2)(x+3) = 0 So, the solution is x = -1, 2, -3

jimthompson5910 (jim_thompson5910):

they just want to factor (not find the roots), so the answer would just be (x+1)(x-2)(x+3)

OpenStudy (anonymous):

nice solve nurali

OpenStudy (anonymous):

\[\lim_{x \rightarrow 2}\frac{ x^{3}+2x^{2}-5x-6 }{ x^{3}-7x+6 }\] was the original question

jimthompson5910 (jim_thompson5910):

factor the denominator and see what cancels then plug in x = 2 and simplify

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

simply solve

OpenStudy (anonymous):

i forget how to do synthetic division?

jimthompson5910 (jim_thompson5910):

this page may help http://www.purplemath.com/modules/synthdiv.htm

OpenStudy (anonymous):

how do i get the divisor for the numerator?

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (anonymous):

for synthetic division

jimthompson5910 (jim_thompson5910):

well the coefficients for the numerator are 1 2 -5 -6 is that what you're asking?

OpenStudy (anonymous):

i got that. what do you put in the bracket in front of those numbers?

jimthompson5910 (jim_thompson5910):

well if you know x+1 is a factor, then -1 is a root, so it would be -1

OpenStudy (anonymous):

thanks